Simplify "(5 + 3i)(5 - 3i)." Is the result real, complex, or both? Explain.
Answer by mathslover(157)

Simplify "(5 + 3i)(5 - 3i)." Is the result real, complex, or both?
Note the expression is of the form (A+B)(A-B)= A^2 - B^2
therefore we can write
5 + 3i)(5 - 3i)
as 5^2 - (3i)^2
=25 - 9i^2
=25 -9(-1) as i^2 = -1
=25+9
=34 which is real number.